It is 8:00 AM in our 24 hour world. What time is it in a 3 hour world? 


1
Expert's answer
2021-11-19T07:05:19-0500

In 24 hour format, there are three shifts of 8 hours that equal to 24 hours (8*3). 8:00 AM in a 24 hour world is one third time of a single day.

"=\frac{8}{24}=\frac{1}{3}"


Then, this time in a 3 hour world is:

"=\frac{1}{3}\times3=1"


In a 3 hour world, 8:00 AM of 24 hour world can be 1 AM or PM.